OpenStudy (anonymous):

suppose in a semester examination 600 student sat for math, 300 physics while 173 student sat for what is the number of student that sat for just a course

OpenStudy (anonymous):

huh? math, physics and [yet to be filled in course] are all courses, is this a trick question? or am I missing something?

OpenStudy (anonymous):

Ah, okay. since the 173 sat in for both math and physics, you want to subtract 173 from both the math and physics students. 600-173 = 427 just math 300-173 = 127 just physics add 427 and 127, and you get the students who took only one test: 554 students

OpenStudy (gw2011):

Since 600 students sat for the math exam and 300 for the physics exam, then a total of 900 students took the exams. Since 173 of these students took both exams, subtract 173 from 900 to get 727 students who took only one exam.

OpenStudy (anonymous):

gw2011 u got it

OpenStudy (anonymous):

I don't think so. the correct answer is 554. Take a simpler example. Suppose only 4 students took math name M1, M2, M3, M4. Now suppose two students take Physics P1 and M1. You see, M1 takes both. So we can list the number of students who took only 1 test: M2 M3 M4 and P1. Mathematically this is: 4 - 1 = 3 + 2 -1 = 1 A total of 4

OpenStudy (gw2011):

If you look at a Venn diagram, you will see that 173 is common to both math and physics. Therefore, the answer of 554 is actually subtracting 173 twice. In other words, if you subtract 173 from math, you are subtracting math and physics from the math part; and if you subtract 173 from physics, you are once again subtracting 173 math and physics from the physics part. The total effect is subtracting 173 twice. Therefore, the 727 should be the correct answer.

OpenStudy (anonymous):

arctan(-1)

OpenStudy (anonymous):

any body to help

OpenStudy (anonymous):

use calculator, press "2ndF" button, followed by "tan", "1" then +/- button at end. dont confuse this with + or - button, it has both + and - on same button. I get -45 degrees.